1.Given utility function U= PxX+PyY    where PX = 12 Birr, Birr, PY = 4 Birr and the income of the consumer is, M= 240 Birr.

A. Find the utility maximizing combinations of X and Y.


B. Calculate marginal rate of substitution of X for Y (MRSX,Y) at equilibrium and interpret your result.


1
Expert's answer
2020-08-02T15:35:32-0400

Solutions:

A. Utility function is given:

U= PxX+PyY

Putting the values of Px and Py in the utility function:

U= 12X + 4Y

Differentiate with respect to X:

"\frac{dU}{dX}= MUx = 12"

Differentiate with respect to Y:

"\frac{dU}{dY}=MUy=4"


At the equilibrium level:

"\frac{MUx}{MUy}=\frac{Px}{Py}"


"\frac{12}{4}=\frac{12}{4}"


"3=3"

Thus, consumer is in equilibrium condition

Budget function:

M = PxX + PyY

240 = 12X + 4Y ......Divide by 4

60 = 3X + Y

Thus, the equilibrium utility maximizing combinations of X and Y is :

3X + Y = 60


B. At the equilibrium level:

MRSxy = "\frac{Px}{Py}"


="\frac{12}{4}"

= 3

Thus, the marginal rate of substitution of X and Y (MRSxy) is 3.

Therefore, it means that the consumer is willing to give up 3 units of X to obtain an additional unit of Y at the same utility level.